OpenStudy (anonymous):

K is the midpoint of LM . The coordinates of K are (5, 12) and the coordinates of L are (2, 6), find the coordinates of M?

OpenStudy (raden):

let the coordinat M is (x,y) so, (2+x)/2 = 5, just solve for x and (6+y)/2 = 12, solve for y

OpenStudy (anonymous):

okee...sory, english not mother language..you understan? i answer you? or is answer?

OpenStudy (raden):

lol, it just a hint. can u do it ?

OpenStudy (anonymous):

i tink so...(: is confused for me..haha

OpenStudy (raden):

well, hope this make sense (2+x)/2 = 5 2+x=2*5 2+x = 10 x = 10 - 2 x = 8

OpenStudy (anonymous):

so y is 18, yes? i so confuse! :(

OpenStudy (raden):

next, (6+y)/2 = 12 6+y = 2*12 6+y = 24 y = 24 - 6 y = 18 ( u are right :d) so, M(8,18)

OpenStudy (anonymous):

okee! tank you much. i no do it if i no had help! sory, i talk french best..

OpenStudy (raden):

nopes... i from indonesia, my english so bad too :p

OpenStudy (anonymous):

oh! see, same! we twins, no? both no talk right! :P

OpenStudy (raden):

hahahaha...

OpenStudy (anonymous):

tank you more...i go now, au revoir twin!

OpenStudy (raden):

welcome :)